# Nonlinear skin modes and fixed-points

### v1

@@ Abstract
We investigate a one-dimensional tight-binding lattice with asymmetrical couplings and various type of nonlinearities to study nonlinear non-Hermitian skin effect. Our focus is on the exploration of nonlinear skin modes through a fixed-point perspective. Nonlinearities are shown to have no impact on the spectral region in the semi-infinite system; however, they induce considerable changes when boundaries are present. The spectrum under open boundary conditions is found not to be a subset of the corresponding spectrum under the semi-infinite boundary conditions. We identify distinctive features of nonlinear skin modes, such as power-energy dependence, degeneracy, and power-energy discontinuity. Furthermore, we demonstrate that a family of localized modes that are neither skin nor scale-free localized modes is formed with the introduction of a coupling impurity. Additionally, we show that an impurity can induce discrete dark and anti-dark solitons.
@@ I Introduction
The non-Hermitian skin effect leads to a large number of localized modes at the boundaries of the lattice, which is not possible in its Hermitian counterpart [1, 2, 3, 4]. This effect can change both the spectral and dynamical properties of a system, giving rise to intriguing dynamical effects such as topological funneling and edge bursts [5, 6, 7]. Scale-free localization is another unique localization phenomena in non-Hermitian systems [8, 9]. This type of localization is different from both skin and Anderson localization as it implies size-dependent localization lengths [10, 11]. Scale-free localization challenges conventional understanding of localization concepts [12, 13, 14, 16, 15, 17, 18, 19]. We note that even a single non-Hermitian defect in an otherwise Hermitian lattice can induce many localized modes whose localization lengths scale with system size [20, 21].
The non-Hermitian skin effect has been extensively studied, but only few papers extending this effect into the nonlinear regime have appeared in the literature [22, 23, 24, 25, 26, 27, 28, 29, 31, 30]. For instance, we explored the nonlinear skin effect and constructed nonlinear skin modes, and discussed fractal structure [22]. M. Ezawa further extended the understanding of this effect by investigating dynamical nonlinear higher-order skin effect and identified a novel topological phase [24]. Experimental evidence of anomalous single-mode lasing, driven by the combined effects of nonlinearity and non-Hermiticity, was provided in [25]. Recent studies have analyzed the nonlinear perturbations of systems with high-order exceptional points, uncovering skin discrete breathers and hierarchical power-law scaling behaviors [26, 27]. In addition to these developments, ref. [28] have explored the emergence of insensitive edge solitons in non-Hermitian topological lattices, which exhibit robustness against perturbations. Reference [29] introduced the concept of solitons with self-induced topological nonreciprocity. An experimental demonstration of the nonlinear non-Hermitian skin effect and skin solitons in an optical system with Kerr nonlinearity has recently been realized [30].
The paper investigates the unique properties and behaviors of nonlinear skin modes in a generalized nonlinear system. We show that fixed points provides us construct nonlinear skin modes and its spectral regions for a finite size lattice and seminfinite lattice. We then outline the key similarities and differences between linear and nonlinear skin modes. Furthermore, we show that a coupling impurity induces a family of localized modes that are neither skin nor scale free localized modes, and also a discrete dark and anti-dark solitons.
@@ II Model
We start with the equation for the field amplitude ${\psi_{n}}$ with $n=1,2,...,N$, describing 1D tight-binding nonlinear lattices with asymmetric couplings $J_{L}$ and $J_{R}$
@@
$(\frac{g_{1}~{}|\psi_{n}|^{2}}{1+\beta~{}|\psi_{n}|^{2}}+g_{2}~{}|\psi_{n}|^{4}+g_{3}~{}\psi_{n}|^{6})~{}\psi_{n}=E~{}\psi_{n}$ (1)
where $\alpha$ and $\beta$ are either 00 or $1$, $g_{1,2,3}$ are all real valued constants, and $E$ is referred as energy for convenience. We set $J_{L}=1$ and $J_{R}=\gamma$ with $10\leq\gamma<$ 1 for simplicity. Eq. (1) specifically describes a system with Kerr-type nonlinearity when $\alpha=\beta=g_{2}=g_{3}=0$, as well as other interesting nonlinearities such as saturable nonlinearity when $\alpha=g_{2}=g_{3}=0$, Ablowitz-Ladik nonlinearity when $\beta=g_{1}=g_{2}=g_{3}=0$, and cubic-quintic-sextic nonlinearity when $\alpha=\beta=0$, among others. We analyze this generalized equation under the following boundary conditions: ${\psi_{N+1}=\psi_{0}=0}$ for open boundary conditions (OBC) and ${\psi_{N+1}=\psi_{1}}$, ${\psi_{0}=\psi_{N}}$ for periodic boundary conditions (PBC), and ${\psi_{\infty}=\psi_{0}=0}$ for the semi-infinite boundary conditions (SIBC). We refer to the SIBC and OBC spectra as the set of values that the parameter $E$ can take for stable skin modes under SIBC and OBC, respectively.
Nonlinear systems exhibit two important features: chaos and fixed points, both of which can be explored in our system using an iterative approach. Starting with arbitrary initial values $\psi_{1}$ and $E$ in Eq. (1), we iteratively compute the subsequent terms $\{\psi_{2},\psi_{3},...,\psi_{N+1}\}$ numerically, and then check if $\psi_{N+1}$ satisfies the required boundary condition. For this reason, we discard solutions where $\psi_{n}$ diverges and consider only bounded solutions. In the latter case, we see either chaotic behavior or fixed points, depending on the values of $\psi_{1}$ and $E$. Chaos arises when a slight variation in $E$ leads to unpredictable behavior as $n$ increases. In contrast, $\psi_{n}$ may converge to a final value, known as the fixed point. This allows us to construct physically acceptable solutions. To find the fixed point, we require that $\psi_{n}$ does not change with $n$ when $n>N_{c}$, i.e., ${\psi_{N_{c}+1}=\psi_{N_{c}+2}=\psi_{N_{c}+3}=a_{0}e^{i\theta}}$ in Eq. (1), where $a_{0}$ is a real number and $\theta$ is an arbitrary parameter. It is worth noting that the critical size $N_{c}$ becomes a finite number at certain values of $\psi_{1}$ and $E$. However, determining these values as a function of the nonlinear parameters is challenging. The equation for $a_{0}$ reads
@@
where $E_{c}=1+\gamma$. This equation yields two types of solutions: the zero fixed point with $a_{0}=0$ and nonzero ones with $a_{0}\neq 0$ being an $E$-dependent real solution of Eq. (2). Note that these fixed points are stable within a specific energy interval. Intriguingly, regardless of the nonlinear interaction strengths, the zero fixed point is stable for the semi-infinite system when $E$ is enclosed by the linear PBC loop in the complex energy plane (defined by $E_{PBC}=e^{ik}+\gamma~{}e^{-ik}$ with $-\pi{\leq}k<\pi$). This can be demonstrated through a linear stability analysis, where a small perturbation is introduced to $\psi_{n}$ in Eq. (1) for $n{>}N_{c}$: ${\psi_{n}\rightarrow a_{0}e^{i\theta}+\epsilon~{}\phi_{n}}$ with $|\epsilon|<<1$ and $\phi_{N_{c}}=\phi_{\infty}=0$. By omitting terms of order $|\epsilon|^{2}$ and higher, we see that the resulting equation does not include any nonlinear parameters when $a_{0}=0$ ($\phi_{n+1}+\gamma~{}\phi_{n+1}=E~{}\phi_{n}$), indicating that $\phi_{n}$ become exactly the linear SIBC eigenstates whose energies are inside the linear PBC loop. Beyond this energy region, it becomes unstable as $\phi_{n}$ diverges. It is interesting to see that the zero fixed point loses its stability in the Hermitian case with $\gamma=1$, as the eigenstates $\phi_{n}$ become extended, causing the zero fixed point to shift to a nonzero value. This explains why nonlinear skin modes are unique to the non-Hermitian system with $\gamma\neq 1$. In this paper, we use these fixed points to construct skin and extended modes, since $a_{0}=0$ and $a_{0}\neq 0$ ensure SIBC (or OBC when $N_{c}$ is finite) and PBC, respectively [33].
@@
The fixed point can also be studied in a bifurcation diagram. For clarity and simplicity, we present a bifurcation diagram for the system with only Kerr-type nonlinearity and real valued $\psi_{1}$ and $E$ (note that $\psi_{1}\rightarrow\psi_{1}e^{i\theta}$ leaves $E$ unaffected and transforms $\psi_{n}$ to $\psi_{n}e^{i\theta}$). Fig. 1(a) shows it for positive real values of $E$ at $\gamma=0.2$ and $g_{1}=1$. The field amplitude always collapses to zero for $E{<}E_{c}=1.20$, and converges to an $E$-dependent nonzero value until $E$ reaches $2.37$, where the nonzero fixed point (period one orbit) becomes unstable, and period-2 points occur, where $\psi_{n}$ starts to alternate between two values instead of converging to a single number. Similarly, at a higher value of $E$, the period-2 points loose stability and bifurcate into period-4 points. This period-doubling behavior continues to period-8, 16 points, etc. until the system exhibits chaotic behavior at $E=2.65$, where a tiny variation in $E$ can lead to a non-periodic behavior in the field amplitude. In this case, the solutions become incompatible with the boundary conditions. The iteration eventually diverges when $E{\geq}2.78$.
@@
Let us now construct extended and skin modes based on the bifurcation diagram in Fig. 1(a). One can obtain extended PBC modes at period-1, period-2, and higher periods until the system enters the chaotic regime, where $\psi_{n}$ no longer converges to a nonzero fixed point. For example, at a period-2 point occurring at $E=2.37$, one can obtain extended modes by setting the initial value $\psi_{1}$ to either of these two values to satisfy PBC. In this case, $\psi_{j}$ oscillates between these two values, causing the corresponding density to exhibit a sawtooth shape, in contrast to the uniform density of the extended modes with a period of $1$. Similarly, at a period-4 point, extended modes oscillate between four distinct values, as shown in the inset of Fig. 1(a). Consider next $E<E_{c}$. In this case, skin modes localized at the left edge appear as $\psi_{n}$ converges to the zero fixed point, satisfying SIBC (and also OBC when $N_{c}$ is finite). As an illustration, we plot the densities of the skin modes up to $n=100$ at $\psi_{1}=0.1$ and $\psi_{1}=0.3$, and two different energies in Fig. 1(b). This plot shows how quickly the field amplitude approaches zero and then gradually decreases to zero, even if $N_{c}$ is infinite. Therefore, such SIBC skin modes may practically be regarded as quasi-stationary OBC modes that maintain their forms over an extended period of time in sufficiently long lattices, as they approximately satisfy OBC [31, 32]. It is worth noting that the localization lengths of the skin modes remain constant with system size but increase with energy $E$, similar to those observed in the linear model. However, the form of the skin modes depends on the power ($P=\sum_{n=1}^{N}|\psi_{n}|^{2}$), as illustrated in the Fig. 1(b), where $\psi_{1}$ is varied from $0.1$ to $0.3$, leading to tighter localization. Note that the power under this change in $\psi_{1}$ increases by a factor of nearly $7$ and $6$ at $E=1$ and $E=1.1$, respectively. It is important to mention that such a scaling of $\psi_{1}$ at fixed $E$ does not generate a distinct skin mode in the linear system. To this end, we note that increasing $\gamma$ extends the skin modes further from the left edge. Our extensive numerical computations across various nonlinear parameters indicate that this extension reaches the lattice size at $\gamma=1$, converting skin modes into extended ones, as illustrated in the inset of Fig. 1(b). This suggests that nonlinear OBC skin modes are unique to the non-Hermitian system, $\gamma\neq 1$, which is in agreement with our previous finding that the zero fixed point loses its stability at $\gamma=1$.
@@
Figure 1: (a) The bifurcation diagram reveals the critical value $E_{c}=1.2$, below which the zero fixed point appears, and above which $E$-dependent fixed points arise. Period doubling occurs at $E=2.37$, followed by chaos at $E=2.65$. In the inset, the densities of the extended PBC modes at a period-4 point with $E=2.6$ are illustrated up to $n=6$. (b) $|\psi_{n}|$ for the skin modes when $\psi_{1}=0.1$ (solid red) and $\psi_{1}=0.3$ (dashed blue) at two different energies at $E=1$ and $E=1.1$ (the more extended ones) are shown. The inset demonstrates that the localization lengths of the skin modes with $\psi_{1}=0.5$ and $E=1$ increase as $\gamma$ is increased from $0.2$ (in black) and $0.7$ (in red) to $0.9$ (in blue), eventually becoming equal to the lattice size at $\gamma=1$ (in green), implying that skin modes are unique to the non-Hermitian systems. (c) The scaled field amplitude as a function of $E$ when $\psi_{1}=0.6$ (in blue) and $\psi_{1}=0.9$ (in orange), and $N=100$. The inset shows the scaled field amplitude at $E=1$ and $N=100$ as a function of $\psi_{1}$, indicating that OBC is satisfied at certain values of $\psi_{1}$. (d) The relation between $E$ and $\psi_{1}$ for OBC modes with $N=100$ in the $E$-interval [0.1,0.2]. The common parameters are $\gamma=0.2$ (except for the inset in (b)), $g=1$ and all other nonlinear interaction strengths are zero.
@@
A portion of SIBC skin modes with finite $N_{c}$ are OBC skin modes, exhibiting novel features such as a continuous spectrum, degeneracies, and energy-power branches. To explore these characteristics, one may begin by iteratively computing $\psi_{N+1}$ as a function of $E$ for an arbitrary value of $\psi_{1}$, and then determine the values of $E$ at which OBC is satisfied, i.e., $\psi_{N+1}(E)=0$. For an illustration, Fig. 1(c) plots the scaled field amplitudes as a function of $E$ at $\psi_{1}=0.6$ and $\psi_{1}=0.9$ when $N=100$. As can be seen, they oscillate between negative and positive values as $E$ increases and touch zero at certain values of $E$. These specific values, where $\psi_{N+1}=0$, correspond to OBC skin modes. As can be seen, selecting a different arbitrary value for $\psi_{1}$ generates distinct value of $E$, so varying $\psi_{1}$ continuously generates a continuum of $E$. Secondly, to illustrate that two or more distinct OBC skin modes can share the same value of $E$ but have different powers, let us fix $E$ and plot $\psi_{N+1}$ as a function of $\psi_{1}$. As shown in the inset of Fig. 1(c), $\psi_{N+1}=0$ at more than one value of $\psi_{1}$, implying degeneracy as distinct OBC skin modes exist at the same value of $E$, each with a different power (correspondingly different $\psi_{1}$). Finally, the modes display an additional characteristic: the energy-power branches. We perform numerical calculations to determine the relation between $E$ and $\psi_{1}$ when ${0.1{\leq}E{\leq}0.2}$. Fig. 1(d) reveals non-smooth behavior between $E$ and $\psi_{1}$. This novel type of branched structure is characterized by dramatic changes in the lowest possible values of $\psi_{1}$ at certain energies, implying that the powers for the OBC skin modes exhibit discontinuities in $E$. For example, an OBC mode is available at $\psi_{1}=0.10$ when $E=0.125$, whereas it appears at $\psi_{1}=0.98$ when $E$ is just below this value.

### v2

@@ Abstract
We investigate a one-dimensional tight-binding lattice with asymmetrical couplings and various type of nonlinearities to study nonlinear non-Hermitian skin effect. Our focus is on the exploration of nonlinear skin modes through a fixed-point perspective. The nonlinear interactions are shown to have no impact on the spectral region in the semi-infinite system; however, they induce considerable changes when boundaries are present. The spectrum under open boundary conditions is found not to be a subset of the corresponding spectrum under the semi-infinite boundary conditions. We identify distinctive features of nonlinear skin modes, such as degeneracy, and power-energy discontinuity. Furthermore, we demonstrate that a family of localized modes that are neither skin nor scale-free localized modes is formed with the introduction of a coupling impurity. Additionally, we show that an impurity can induce discrete dark and anti-dark solitons.
@@ I Introduction
The non-Hermitian skin effect leads to a large number of localized modes at the boundaries of the lattice, which is not possible in its Hermitian counterpart [1, 2, 3, 4, 5]. This effect can change both the spectral and dynamical properties of a system, giving rise to intriguing dynamical effects such as topological funneling and edge bursts [6, 7, 8]. Scale-free localization is another unique localization phenomena in non-Hermitian systems [9, 10]. This type of localization is different from both skin and Anderson localization as it implies size-dependent localization lengths [11, 12]. Scale-free localization challenges conventional understanding of localization concepts [13, 14, 15, 17, 16, 18, 19, 20, 21]. We note that even a single non-Hermitian defect in an otherwise Hermitian lattice can induce many localized modes whose localization lengths scale with system size [22, 23].
The non-Hermitian skin effect has been extensively studied, but only few papers extending this effect into the nonlinear regime have appeared in the literature [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. For instance, we explored the nonlinear skin effect and constructed nonlinear skin modes, and discussed fractal structure [24]. M. Ezawa further extended the understanding of this effect by investigating dynamical nonlinear higher-order skin effect and identified a novel topological phase [26]. Experimental evidence of anomalous single-mode lasing, driven by the combined effects of nonlinearity and non-Hermiticity, was provided in [27]. Recent studies have analyzed the nonlinear perturbations of systems with high-order exceptional points, uncovering skin discrete breathers and hierarchical power-law scaling behaviors [28, 29]. In addition to these developments, ref. [30] have explored the emergence of insensitive edge solitons in non-Hermitian topological lattices, which exhibit robustness against perturbations. Reference [31] introduced the concept of solitons with self-induced topological nonreciprocity. An experimental demonstration of the nonlinear non-Hermitian skin effect and skin solitons in an optical system with Kerr nonlinearity has recently been realized [32].
The paper investigates the unique properties and behaviors of nonlinear skin modes in a generalized nonlinear system by using the numerical shooting method. We show that fixed points provides us construct nonlinear skin modes and its spectral regions for a finite size lattice and semi-infinite lattice. We then outline the key similarities and differences between linear and nonlinear skin modes. Furthermore, we show that a coupling impurity induces a family of localized modes that are neither skin nor scale free localized modes, and also a discrete dark and anti-dark solitons.
@@ II Model
We start with the equation for the complex field amplitude ${\psi_{n}}$ with $n=1,2,...,N$, describing one-dimensional tight-binding nonlinear lattices with asymmetric couplings $J_{L}$ and $J_{R}$
@@
$(\frac{g_{1}~{}|\psi_{n}|^{2}}{1+\beta~{}|\psi_{n}|^{2}}+g_{2}~{}|\psi_{n}|^{4}+g_{3}~{}|\psi_{n}|^{6})~{}\psi_{n}=E~{}\psi_{n}$ (1)
where $\alpha$ and $\beta$ are either 00 or $1$, $g_{1,2,3}$ are all real valued constants, and $E$ is referred as energy for convenience. We set ${J_{L}=1}$ and ${J_{R}=\gamma}$ with $10\leq\gamma<$ 1 for simplicity. This equation contains five nonlinear parameters that can be chosen independently. It specifically describes a system with Kerr-type nonlinearity when ${\alpha=\beta=g_{2}=g_{3}=0}$, as well as other interesting nonlinearities such as saturable nonlinearity when ${\alpha=g_{2}=g_{3}=0}$ [36], Ablowitz-Ladik nonlinearity when ${\beta=g_{1}=g_{2}=g_{3}=0}$ [37], and cubic-quintic-sextic nonlinearity when ${\alpha=\beta=0}$ [38], among others. We analyze this generalized equation under the following boundary conditions: ${\psi_{0}=\psi_{N+1}=0}$ for open boundary conditions (OBC) and ${\psi_{1}=\psi_{N+1}}$, ${\psi_{0}=\psi_{N}}$ for periodic boundary conditions (PBC), and ${\psi_{0}=\psi_{\infty}=0}$ for the semi-infinite boundary conditions (SIBC). We refer to the SIBC and OBC spectra as the set of values that the parameter $E$ can take for stable skin modes localized at the left edge under SIBC and OBC, respectively.
Nonlinear systems exhibit two important features: chaos and fixed points. They can be explored in our system using the shooting method, which converts the boundary value problem into an initial value problem by guessing the field amplitude at the left edge. In this way, we obtain a unique solution that satisfies the boundary condition. Starting with an arbitrary initial value $\psi_{1}$ and $E$, we perform numerical iterations in Eq.(II) to compute the subsequent terms ${\{\psi_{2},\psi_{3},...,\psi_{N+1}\}}$, and then check if the remaining boundary condition for ${\psi_{N+1}}$ is satisfied. If it is not, the initial guess $\psi_{1}$ and $E$ are modified, and the numerical iteration is repeated until the solution meets the required condition for ${\psi_{N+1}}$. For bounded solutions (non-diverging $\psi_{n}$), our numerical analysis reveal either chaotic behavior or fixed points at high iteration values, depending on $E$. Chaos arises when a slight variation in $E$ in the numerical iterations leads to unpredictable behavior as $n$ increases. In contrast, $\psi_{n}$ may converge to a final value, known as a fixed point (also known as an equilibrium point). A fixed point is a point in the system where $\psi_{n}$ does not change with $n$ when ${n>N_{c}}$, where $N_{c}$ is the critical size. Furthermore, this fixed point is stable if small perturbations around the fixed point decay with $n$, causing the system to converge back to the fixed point.
To find the fixed points in our system, we require that $\psi_{n}$ remain unchanged when $n>N_{c}$, i.e., ${\psi_{N_{c}+1}=\psi_{N_{c}+2}=...=\psi_{N}=a_{0}e^{i\theta}}$ in Eq.(II), where $a_{0}$ and $\theta$ are real numbers. It is worth noting that the critical size $N_{c}$ is a finite number at certain values of $\psi_{1}$ and $E$. However, determining these values as a function of the nonlinear parameters $\alpha$, $\beta$ and ${g_{1,2,3}}$ is challenging. The equation for $a_{0}$ reads
@@
where $E_{c}=1+\gamma$. It yields two types of solutions: the zero fixed point with ${a_{0}=0}$ and nonzero ones with ${a_{0}\neq 0}$, which are real-valued solutions of this equation. Unlike the zero fixed point, the nonzero one is $E$-dependent and appears when ${E>E_{c}}$ (since ${a_{0}^{2}>0}$). For example, ${a_{0}^{2}=\frac{E-E_{c}}{E_{c}}}$ in the case of the Ablowitz-Ladik nonlinearity, i.e., ${g_{1,2,3}=0}$ and ${\alpha=1}$. Eq.(2) also reveals that $\theta$ is an arbitrary parameter. In the numerical shooting method, changing the initial value ${\psi_{1}\rightarrow\psi_{1}e^{i\theta}}$ leaves $E$ unaffected and transforms the remaining $\psi_{n}$ to ${\psi_{n}e^{i\theta}}$. Therefore, for the sake of simplicity, we consider only real values of $\psi_{1}$ throughout this work.
@@
Let us next study the stability of the fixed points. Intriguingly, the stability of the zero fixed point for the semi-infinite system is controlled by the non-Hermiticity rather than the nonlinear interactions. To be precise, the zero fixed point is stable regardless of the nonlinear interaction strengths for the semi-infinite system when $E$ is enclosed by the linear PBC loop in the complex energy plane (defined by $E_{PBC}=e^{ik}+\gamma~{}e^{-ik}$ with $-\pi{\leq}k<\pi$). This can be demonstrated through a linear stability analysis, where a small perturbation is introduced to $\psi_{n}$ in Eq.(II) for $n{>}N_{c}$: ${\psi_{n}\rightarrow a_{0}e^{i\theta}+\epsilon~{}\phi_{n}}$ with $|\epsilon|<<1$ and ${\phi_{\infty}=0}$. By omitting terms of order $|\epsilon|^{2}$ and higher, we see that the resulting equation does not include any nonlinear parameters when $a_{0}=0$: ${\phi_{j+1}+\gamma~{}\phi_{j-1}=E~{}\phi_{j}}$ with ${j=n+N_{c}}$. The condition ${\phi_{\infty}=0}$ is satisfied when $E$ is inside the linear PBC loop in the complex energy plane. Beyond this energy region, the zero fixed point becomes unstable as a small perturbation does not converge to zero. This analysis also reveals another interesting result: The zero fixed point would lose its stability if the system is Hermitian $\gamma=1$, as there exists extended eigenstates $\phi_{j}$ in the energy interval ${[-E_{c},E_{c}]}$. This explains why the nonlinear skin modes are unique to the non-Hermitian system with $\gamma\neq 1$. Unlike the zero fixed point, the stability region of the nonzero fixed point strongly depends on the nonlinear parameters. They are typically stable within a narrow range of energy. Fortunately, they can be stable across an extended range of energy values in some particular cases. For example, we repeat the stability analysis for only the Ablowitz-Ladik nonlinearity, where ${a_{0}^{2}=\frac{E-E_{c}}{E_{c}}}$. Assuming real valued function $\phi_{j}$ with real energy ${E}$, we obtain ${\phi_{j+1}+\gamma~{}\phi_{j-1}=E^{\prime}~{}\phi_{j}}$, where ${E^{\prime}=E_{c}(\frac{2E_{c}}{E}-1)}$. When ${E>E_{c}}$, we see that ${E_{c}<E^{\prime}<E_{c}}$, implying converging solutions, $\phi_{\infty}=0$. This proves the stability of the nonzero fixed point in the presence of only the Ablowitz-Ladik nonlinearity. In this paper, we consider these fixed points to construct skin and extended modes using the simple shooting method. Specifically, $a_{0}=0$ and $a_{0}\neq 0$ ensure SIBC (or OBC when $N_{c}$ is finite) and PBC, respectively [40].
@@
We can study bifurcation diagrams for real positive values of $E$, where $E$ is treated as the bifurcation parameter. From the earlier discussion, one can get a general understanding of the bifurcation diagrams as we increase $E$ from zero. In such diagrams, $\psi_{n}$ converges to zero value regardless of the nonlinear parameters when ${E<E_{c}}$. Above this point, the zero fixed point becomes unstable, and $\psi_{n}$ converges to an $E$-dependent nonzero fixed value. Then, the nonzero fixed point becomes unstable at a certain value of $E$ that strongly depends on the nonlinear interaction strengths. This leads to bifurcations and ultimately to chaos (bifurcation does not appear if the nonzero fixed point is stable). As a result, in these diagrams, changing the nonlinear parameters at fixed $\gamma$ changes the points where bifurcation and chaos occur, while leaving $E_{c}$ unchanged. With these in mind, one can explore bifurcation diagrams with various values of the nonlinear parameters. Due to its experimental relevance, we present a bifurcation diagram for the system with only Kerr-type nonlinearity. Fig. 1(a) shows it for positive values of $E$ at ${\gamma=0.2}$ and ${g_{1}=1}$. The field amplitude always collapses to zero for ${E{<}E_{c}=1.20}$, and converges to an $E$-dependent nonzero value until $E$ reaches $2.37$, where the nonzero fixed point (period one orbit) becomes unstable, and period-2 points occur, where $\psi_{n}$ starts to alternate between two values instead of converging to a single number. Similarly, at a higher value of $E$, the period-2 points loose stability and bifurcate into period-4 points. This period-doubling behavior continues to period-8, period-16, and higher points, until the system exhibits chaotic behavior at ${E=2.65}$, where a tiny variation in $E$ can lead to a non-periodic behavior in the field amplitude $\psi_{n}$. In this case, the solutions become incompatible with the boundary conditions. The iteration eventually diverges when ${E{\geq}2.78}$.
@@
Let us now construct extended and skin modes based on the bifurcation diagram in Fig. 1(a). One can obtain extended PBC modes at period-1, period-2, and higher periods until the system enters the chaotic regime. For example, at a period-2 point occurring at $E=2.37$, $\psi_{j}$ oscillates between two values, causing the corresponding density to exhibit a sawtooth shape, in contrast to the uniform density of the extended modes with a period of $1$. Similarly, at a period-4 point, extended modes oscillate between four distinct values, as shown in the inset of Fig. 1(a). Note that we set the initial value $\psi_{1}$ to either of these four values in the numerical shooting method. Consider next ${E<E_{c}}$ for the system under SIBC or OBC. In this case, skin modes localized at the left edge appear as $\psi_{n}$ converges to the zero fixed point, satisfying SIBC (and also OBC when $N_{c}$ is finite). As an illustration, we perform iteration for two different initial values of $\psi_{1}$ and plot the densities of the skin modes up to ${n=100}$ at ${E=1.0}$ (black) and ${E=1.1}$ (red) in Fig. 1(b). Note that these modes are tightly localized at the left edge even if $N_{c}$ is infinite. Therefore, such SIBC skin modes may practically be regarded as quasi-stationary OBC modes that maintain their forms over an extended period of time in sufficiently long lattices, as they approximately satisfy OBC [35, 39]. As can be seen from the figure, the localization lengths of these skin modes increase with energy $E$ but remain constant with system size $N$, similar to those observed in the linear model. However, the localization lengths depend on the power ($P=\sum_{n=1}^{N}|\psi_{n}|^{2}$), as illustrated in the figure, where varying $\psi_{1}$ from 0.10.1{0.1}0.1 (solid) to 0.30.3{0.3}0.3 (dashed) increases the power by a factor of nearly $7$ and $6$ for ${E=1.0}$ and ${E=1.1}$, respectively. Note that such a scaling of $\psi_{1}$ at fixed $E$ does not generate a distinct skin mode in the linear system. The inset of Fig. 1(b) shows the densities of the skin modes at ${\psi_{1}=0.5}$ and ${E=1.0}$ at four different values of $\gamma$: $0.2$ (in black), $0.7$ (in red), 0.90.9{0.9}0.9 (in blue), and 1.01.0{1.0}1.0 (in green). Increasing $\gamma$ extends the skin modes further from the left edge, as reduced asymmetry in the couplings weakens skin localization, leading to more extended skin modes. Notably, this extension reaches the lattice size at ${\gamma=1}$, converting skin modes into extended ones. This suggests that nonlinear skin modes are unique to the non-Hermitian system, ${\gamma\neq 1}$, which is also in agreement with our previous finding that the zero fixed point loses its stability at ${\gamma=1}$, implying the absence of skin modes.
@@
Figure 1: (a) The bifurcation diagram reveals the critical value ${E_{c}=1.2}$, below which the zero fixed point appears, and above which $E$-dependent fixed points arise. Period doubling occurs at ${E=2.37}$, followed by chaos at ${E=2.65}$. In the inset, the densities of the extended PBC modes at a period-4 point with ${E=2.6}$ are illustrated up to $n=6$. (b) $|\psi_{n}|$ for the skin modes when ${\psi_{1}=0.1}$ (solid) and ${\psi_{1}=0.3}$ (dashed) at two different energies at $E=1.0$ (black) and ${E=1.1}$ (red). The inset demonstrates $|\psi_{n}|$ at ${\psi_{1}=0.5}$ and ${E=1.0}$ as $\gamma$ is increased from 0.20.2{0.2}0.2 (in black) and 0.70.7{0.7}0.7 (in red) to ${\gamma=0.9}$ (in blue) and 1.01.0{1.0}1.0 (in green). (c) The scaled field amplitude as a function of $E$ at ${\psi_{1}=0.6}$ (in blue) and ${\psi_{1}=0.9}$ (in orange), and $N=100$. The inset shows the scaled field amplitude at ${E=1.0}$ and ${N=100}$ as a function of $\psi_{1}$, indicating that OBC is satisfied at certain values of $\psi_{1}$. (d) The relation between $E$ and $\psi_{1}$ for OBC modes with $N=100$ in the $E$-interval [0.1,0.2]. The common parameters are ${\gamma=0.2}$ (except for the inset in (b)), ${g_{1}=1}$ and all other nonlinear interaction strengths are zero.
@@
A portion of SIBC skin modes with finite $N_{c}$ are OBC skin modes. These modes exhibit some features, such as continuous spectrum, degeneracies, and energy-power branches. To explore these characteristics, one may begin by iteratively computing $\psi_{N+1}$ as a function of $E$ for an arbitrary value of $\psi_{1}$, and then determine the values of $E$ at which OBC is satisfied, i.e., $\psi_{N+1}(E)=0$. As an illustration, we consider Kerr-type nonlinearity with ${g_{1}=1}$ for a finite lattice with $N=100$ and ${\gamma=0.2}$. Fig. 1(c) plots the scaled $\psi_{N+1}$ as a function of $E$ at ${\psi_{1}=0.6}$ (in blue) and ${\psi_{1}=0.9}$ (in orange). As can be seen, they oscillate as $E$ varies and touch zero at certain values of $E$. These specific values, where $\psi_{N+1}=0$, correspond to OBC skin modes at $N=100$. This plot also shows that selecting a different arbitrary value for $\psi_{1}$ generates distinct values of $E$, so varying $\psi_{1}$ continuously generates a continuum of $E$. Note that these solutions can also be used to construct the OBC skin modes for larger lattice sizes, as $\psi_{N+1}=0$ guarantees that $\psi_{n>N+1}=0$ (a zero fixed point). Our numerical calculations reveal that $N_{c}$ does not take finite values when $E$ is complex, suggesting that the nonlinear OBC spectrum is real valued, similar to the linear OBC spectrum. In addition to the continuous spectrum, OBC skin modes exhibit another feature: two or more distinct OBC skin modes with different powers can share the same value of $E$. To illustrate this, we plot $\psi_{N+1}$ as a function of $\psi_{1}$ at ${E=1}$ in the inset of Fig. 1(c). As can be seen, the condition ${\psi_{N+1}=0}$ is satisfied at multiple values of $\psi_{1}$, implying degeneracy, as distinct OBC skin modes exist at the same value of $E$, each with a different $\psi_{1}$ (correspondingly a different power). Finally, the OBC modes display an additional feature: the energy-power branches. We perform numerical calculations to determine the relation between $E$ and $\psi_{1}$ when ${0.1{\leq}E{\leq}0.2}$. Fig. 1(d) reveals non-smooth behavior between $E$ and $\psi_{1}$. This novel type of branched structure is characterized by dramatic changes in the lowest possible values of $\psi_{1}$ at certain energies, implying that the powers for the OBC skin modes exhibit discontinuities in $E$. For example, an OBC mode is available at $\psi_{1}=0.10$ when $E=0.125$, whereas it appears at $\psi_{1}=0.98$ when $E$ is just below this value.